OpenStudy (anonymous):

Find the slope of the line whose equation is 4y - 3x + 6 = 0. -3/4 -3/2 3/4

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

4y-3x+6=0 4y-3x=6 4y=3x+6 y=3/4x + 6 <- slope intersept form